The mean and variance from a group of 80 observations are 63.2 and 25.93 respectively. If 60 of these observations have nean 64.8 and SD=4 find mean and SD of remaining 20

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Akash
For mean
Total = 80 * 63.2 = 5056
Sum of 60 = 60* 64.8 = 3888
Sum of remaining 20 = 5056-3888 = 1168
Hence
Mean of remaining 20 = 1168/20 = 58.4
 
Regards
Arun (askIITians forum expert)